The root-mean-square speed of molecules of
mass m in an ideal gas at temperature T is

please give process

Grade:9

1 Answers

Sunil Kumar FP
10 years ago
we know that the kinetic energy of a gas molecule is given by E=1/2mv^2
v is the root mean speed.
but it is equal to 3/2kT where k is boltzman constant,Tis the temperature.
1/2mv^2=3/2kT
v=under root 3kT/m ,m is the mass of each gas molecule

thanks and regards
sunil kr
askIItian faculty